OpenStudy (anonymous):

Fidn a quadratic equation with roots -1+4i and -1-4i

OpenStudy (anonymous):

Find*

OpenStudy (anonymous):

(x - root)(x - root), so you need to mulitply \[(x + 1 - 4i)(x+1+4i)\] When you do the imaginary parts will cancel and you'll have the quadratic

OpenStudy (anonymous):

@peachpi is right

OpenStudy (anonymous):

I still don't understand it very well... i'm sorry @peachpi

OpenStudy (anonymous):

do you know how to foil?

OpenStudy (anonymous):

No not very well @peachpi

OpenStudy (anonymous):

that's ok. basically it's just applying the distributive property. You take each term from the 1st parentheses and multiply it by each term in the second. Starting with the \(x\) in the 1st, multiply and get \(x(x+1+4i)=x^2+x+4ix\) You try it with the 2nd and third terms. Just distribute. 2nd: \(1(x+1+4i)=\) 3rd: \(-4i(x+1+4i)=\)

OpenStudy (anonymous):

2nd: x+1+4i 3rd: -4ix-4i-16i right? @peachpi

OpenStudy (anonymous):

The second one is right. The third is almost right. -16i should be -16i². Then because \(i^2=-1\), it reduces to 16. Altogether is should be \(-4ix-4i+16\) Make sense?

OpenStudy (anonymous):

Yep makes sense now! @peachpi

OpenStudy (anonymous):

so now all you have to do is add up what we got from the multiplication for each step. Every term with an i will cancel, and then combine everything else to get the final answer

OpenStudy (anonymous):

Okay thanks so much @peachpi